OpenStudy (anonymous):

An aquarium wants to install a new fish tank with a parabolic glass dome fitted at its bottom to allow visitors to look at the marine life from below. The glass dome will be 60 feet in diameter and will have a depth of 4 feet. Find the equation of the parabolic cross-section of the dome. y2 = 16/30(x) y2 = 4/15(x) y = 1/225(x^2) y = 15/14(x^2) y = 1/15(x^2)

OpenStudy (amistre64):

well, if we want to match the directions exactly, we want to move the 'parent' parabola around such that: y = ax^2 + b matches our design 60/2 = 30 so we want y=0 when x=30, and y=-4 when x=0 -4 = a0^2 + b , can solve for b; and then knowing b we can solve for a

OpenStudy (amistre64):